Find the slope-intercept form of the equation of the line passing through the given points.
(2, 2),(4, −3/2)

Answer by Cromlix(4381)

Using equation for slope:
y2 - y1/x2 - x1
-3/2 - 2/4 - 2
-7/2/2
= -7/4 (slope)
Using y = mx + c with coords (2,2)
2 = -7/4 * 2 + c
2 = -14/4 + c
2 + 14/4 = + c
22/4 = c
y = -7/4x + 22/4
Multiply thro' by 4
4y = -7x + 22
Hope this helps.
:-)